OpenStudy (he66666):

Find a 2x1 matrix x with entries not all zero such that Ax=3x, where A = [2 1; 1 2]. Am I supposed to use the Gaussian elimination method first? I tried it but I can't make matrix A into row echelon form.. can someone please help me?

OpenStudy (kinggeorge):

Just let your matrix \[x=\left[\begin{matrix}x_1\\x_2\end{matrix}\right]\] and multiply your two matrices. Once you do that, solve for the values so that you get 3x.

OpenStudy (kinggeorge):

Your solution should be something like \[Ax=\left[\begin{matrix} 2x_1+x_2 \\x_1+2x_2\end{matrix}\right]\]And you want \[2x_1+x_2=3x_1\]\[x_1+2x_2=3x_2\]With this you can do some quick substitutions and get an answer.

OpenStudy (he66666):

Thanks a lot! Is this the only way to solve this though? How I tried was subtract both sides of Ax=3x by 3x so you get Ax-3x=0 which becomes (A-3I)x=0. But then it didn't work..

OpenStudy (kinggeorge):

I think your mistake in that was saying that 3x=3Ix. To have the identity matrix, you need it to be square, and x is not a square matrix.

OpenStudy (anonymous):

what he has should still lead to a solution. Once you have\[(A-3I)x=0\]You should realize you are looking for a basis for the Null Space (or Kernel) of the matrix A-3I

OpenStudy (he66666):

Oh I see. So for similar questions like these, you would use the same techniques? I have another question that's like: Find a 3x1 matrix x with entries not all zero such that Ax=1x, where A = [1 2 -1; 1 0 1; 4 -4 5]. So I would set up x = [x1, x2, x3]?

OpenStudy (anonymous):

he66666, do you know about eigenvalues or eigenvectors yet?

OpenStudy (he66666):

@joemath314159 We haven't learnd Null space yet so I don't really know..

OpenStudy (he66666):

No, we didn't learn about those either. We just started learning solving linear systems using the Gaussian elimination method.

OpenStudy (anonymous):

wow. Your professor is very sneaky then lol. Since gaussian elimination is all you have, KingGeorge's method is best.

OpenStudy (anonymous):

Once you learn things like null space, eigenvalues and whatnot, you should come back to this problem to have a good laugh.

OpenStudy (he66666):

Oh haha. I guess I'll use this technique to solve these problems. Thanks for your help joemath314159 and KingGeorge :)

OpenStudy (kinggeorge):

You're welcome.